Grade 11General Physics

Calculate the number of days in one year when the average distance between the sun and earth becomes half to its initial value

Sunil vyas
8 Years agoGrade 11

1 Answer

Khimraj
8 Years ago
According to Kepler’s law of time Periods, the square of time period of a planet is directly proportional to cube of the semimajor axis of the orbit.Let T1 and a1 are initial time period and distance between sun and earth.So(T12/a13) = (T22/a23)(T1/T2)2 = (a1/a2)3(T1/T2)2 = (2)3(T1/T2)2 = 8T2 = T1/(81/2)T2 = 365/(81/2) = 129 Days.Hope it clears if still you have doubt you can ask me.If you like answer then please approve it.